# Quantum Natural Gradient with Geodesic Corrections for Small Shallow Quantum Circuits

### v2

@@ 5 Computing the Christoffel Symbols with the Parameter-Shift Rule
$g_{j_{1}j_{2}}(\bm{\theta})=-\frac{1}{2}\frac{\partial^{2}}{\partial\theta_{j_{1}}\partial\theta_{j_{2}}}\left.\left|\langle\psi(\bm{\theta}^{\prime})|\psi(\bm{\theta})\rangle\right|^{2}\right|_{\theta^{\prime}=\theta}$ (24)
Using the parameter-shift rule, this metric can be reformulated as:
@@
where $\mathbf{e}_{j_{1}}$ and $\mathbf{e}_{j_{2}}$ are unit vectors in the parameter space, and the shifted parameter is given by $s=\pi/2$.
If the metric tensor simplifies to diagonal elements, we have:
@@
In this scenario, the Christoffel symbols can be directly computed from the derivatives of the metric tensor using the higher-order parameter-shift rule, enabling their extraction directly from the quantum circuit. This approach bypasses traditional classical computations, allowing for efficient and accurate estimation directly on the quantum hardware. For the diagonal case, the results can be encapsulated in the following proposition:
@@
$\Gamma^{i}_{jk}$ $=\frac{1}{2g_{ii}(\theta)}\Bigg{[}\frac{1}{8}\delta_{ij}\Bigg{(}-\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}+\dfrac{\pi}{2}e_{k})\rangle\right|^{2}+\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}-\dfrac{\pi}{2}e_{k})\rangle\right|^{2}\Bigg{)}$ (27)
$+\frac{1}{8}\delta_{ik}\Bigg{(}-\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}+\dfrac{\pi}{2}e_{j})\rangle\right|^{2}+\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}-\dfrac{\pi}{2}e_{j})\rangle\right|^{2}\Bigg{)}$
$-\frac{1}{8}\delta_{jk}\Bigg{(}-\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{j}+\dfrac{\pi}{2}e_{i})\rangle\right|^{2}+\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{j}-\dfrac{\pi}{2}e_{i})\rangle\right|^{2}\Bigg{)}\Bigg{]}$
@@
6 Application Examples
@@
6.1 Example 1 : Single Qubit
@@
The Hamiltonian for which we want to find the ground state is $H=\sigma_{x}$. Thus, the cost function for the system is:
@@
The Christoffel symbol of this metric, calculated using equation (4), the no zeros terms are:
@@
Figure 1: Performance of GD, QNG, and QNGGC with varying correction parameters ($b=\eta^{2}$, $b=0.05$, and $b=0.2$). The initial parameters are set to $[\pi/12,\pi/12]$, and a learning rate $\eta=0.05$ is used for all optimizers. (a) Shows the evolution of the energy cost function over 30 iterations, demonstrating the improved convergence behavior of QNGGC relative to QNG and GD. (b) Depicts the logarithmic energy difference, $\log_{10}(\Delta E)$, where $\Delta E$ represents the gap between the current energy and the target energy $E^{*}=-1$. The steeper slopes of QNGGC illustrate its enhanced convergence rate. (c) Displays the norm of the geodesic correction term across different correction parameters $b$, highlighting the impact of geodesic corrections during optimization. (d) Shows the trajectory of the parameters $\theta_{0}$ and $\theta_{1}$ throughout the optimization process, illustrating how each method navigates the parameter space. The QNGGC paths are more direct, emphasizing the benefits of incorporating geodesic corrections.
Figure 1 demonstrates the impact of geodesic corrections on optimization performance. The inclusion of geodesic corrections in QNGGC significantly enhances convergence speed (in terms of steps) compared to GD and QNG, as depicted in subplots (a) and (b), where QNGGC rapidly approaches the target energy $E^{*}=-1$. Subplot (c) shows the norm of the geodesic correction term, $\frac{b}{2}\Gamma^{i}_{lm}(\bm{\theta}_{t})\left(g^{lj}(\bm{\theta}_{t})\partial_{j}\mathcal{L}(\bm{\theta}_{t})\right)\left(g^{mj}(\bm{\theta}_{t})\partial_{j}\mathcal{L}(\bm{\theta}_{t})\right)$, which plays a crucial role in guiding the optimization process, particularly in the initial iterations. The subsequent reduction of this norm towards zero indicates that the convergence of the gradient is being effectively managed, leading to more stable updates. Subplot (d) illustrates the optimization paths in the parameter space, where QNGGC provides a more direct and efficient trajectory compared to GD and QNG. This behavior underscores the advantage of leveraging curvature corrections, resulting in more precise and adaptive parameter updates.
6.2 Example 2: Two-Qubit Simulation of the Hydrogen Molecule ($\text{H}_{2}$)
@@
where $\alpha=0.4$ and $\beta=0.2$. In this Hamiltonian, $\sigma_{z}$ and $\sigma_{x}$ are Pauli matrices that represent spin operators acting on individual qubits. This Hamiltonian has four eigenvalues:
@@
In this example, the Fubini-Study metric $F$ is calculated as:
@@
Figure 3: Comparison of the convergence of GD, QNG, and QNGGC. Initial parameters are $[-0.2,-0.2,0]$, with a learning rate $\eta=0.05$ for all optimizers and $b=\eta^{2}$ for QNGGC. Subplot (a) shows energy convergence, (b) plots $\log_{10}(\Delta E)$, where $\Delta E$ is the energy difference from the target $E^{*}\approx-0.82462$, (c) shows the norm of the geodesic correction term in QNGGC, and (d) illustrates the fidelity convergence rate. The optimization runs for 30 iterations with a regularization term $\lambda=10^{-6}$.
@@
Figure 4: Same conditions as in Figure 3, except the initial parameters are set to $[\pi/2,\pi/2,0]$ and $b=0.4$.
Figure 4 uses the same conditions as Figure 3, but with a different value of $b=0.4$ and initial parameters set to a more challenging configuration, $[\pi/2,\pi/2,0]$. This demonstrates how the QNGGC optimizer efficiently identifies a path to escape from plateaus, leading to faster convergence to the ground state compared to GD and QNG.
@@
6.3 Example 3: Transverse Field Ising Model
@@
where $Z_{i}$ and $X_{i}$ are Pauli operators acting on the $i$-th qubit, with the transverse field strength fixed at $h=10$ for our simulations. This model captures the interplay between nearest-neighbor spin interactions and a transverse magnetic field, driving the system between distinct quantum phases. For smaller values of the transverse field $h$, the spins tend to align along the z-direction, corresponding to a ferromagnetic ordered phase. As $h$ increases, the system becomes increasingly disordered and transitions into a paramagnetic phase. The critical point at $h=1$ marks the quantum phase transition between these phases, where spontaneous symmetry breaking and critical quantum fluctuations occur.
@@ 7 Conclusions and Outlook
The QNGGC update rule reduces the number of steps needed to reach the target energy but remains time-intensive compared to native QNG due to the computational cost of Christoffel symbol calculations. Currently, QNGGC is mainly applicable to small, shallow quantum circuits.
Future work could explore extending our results to noisy and nonunitary circuits, following approaches such as those presented in [10]. Developing more efficient methods for computing Christoffel symbols, such as stochastic parameter-shift rules and simultaneous perturbation stochastic approximation techniques for the Quantum Fisher Information [11], would also be valuable. Additionally, randomness-based methods for Christoffel symbol computation [12] represent a promising avenue for further research. These extensions would enhance the broader applicability and efficiency of geodesic corrections in quantum optimization frameworks. Further research will also focus on moving beyond second-order approximations. Our approach could also be applied to time-dependent optimizers, particularly those inspired by the metric tensor and differential geometry.
8 Acknowledgements
@@ A.2 Fubini-Study Metric
To derive the Fubini-Study metric tensor $g_{ij}(\bm{\theta})$, we begin by examining the infinitesimal distance between nearby quantum states in the parameter space. The metric tensor captures the local geometric structure of this space, reflecting how the quantum state $|\psi(\bm{\theta})\rangle$ evolves under small displacements [6] [8] [32].
@@
Differentiating this condition twice with respect to the parameters $\theta^{i}$ and $\theta^{j}$, we obtain:
$\langle\psi(\bm{\theta})|\partial_{i}\partial_{j}\psi(\bm{\theta})\rangle+\langle\partial_{i}\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\rangle+\langle\partial_{j}\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\rangle=0.$ (42)
Next, consider the quantum state $|\psi(\bm{\theta}+\delta\theta)\rangle$ near $|\psi(\bm{\theta})\rangle$, where $\delta\theta$ is the displacement vector. Using a Taylor expansion, we express the shifted state as:
@@
$\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\delta\theta)\rangle=1+\langle\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\rangle\delta\theta^{i}+\frac{1}{2}\langle\psi(\bm{\theta})|\partial_{i}\partial_{j}\psi(\bm{\theta})\rangle\delta\theta^{i}\delta\theta^{j}+\mathcal{O}(\delta\theta^{3}).$ (44)
The fidelity between the states $|\psi(\bm{\theta})\rangle$ and $|\psi(\bm{\theta}+\delta\theta)\rangle$ is given by:
$|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\delta\theta)\rangle|^{2}$ $=1+\langle\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\rangle\delta\theta^{i}+\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\rangle\delta\theta^{i}$
$\quad+\frac{1}{2}\langle\psi(\bm{\theta})|\partial_{i}\partial_{j}\psi(\bm{\theta})\rangle\delta\theta^{i}\delta\theta^{j}+\mathcal{O}(\delta\theta^{3}).$ (45)
Using the normalization and differentiation results, we find the infinitesimal squared distance between the quantum states as:
$d^{2}(P_{\psi},P_{\psi+\delta\theta})=\text{Re}\left[\langle\partial_{i}\psi|\partial_{j}\psi\rangle-\langle\partial_{i}\psi|\psi\rangle\langle\psi|\partial_{j}\psi\rangle\right]\delta\theta^{i}\delta\theta^{j}.$ (46)
This form directly defines the Fubini-Study metric tensor:
$g_{ij}(\bm{\theta})=\text{Re}\left[\langle\partial_{i}\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\rangle-\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\rangle\langle\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\rangle\right].$ (47)

### v3

@@ V Computing the Christoffel Symbols with the Parameter-Shift Rule
$g_{j_{1}j_{2}}(\bm{\theta})=-\frac{1}{2}\frac{\partial^{2}}{\partial\theta_{j_{1}}\partial\theta_{j_{2}}}\left.\left|\langle\psi(\bm{\theta}^{\prime})|\psi(\bm{\theta})\rangle\right|^{2}\right|_{\bm{\theta}^{\prime}=\bm{\theta}}$ (26)
To evaluate this second derivative, we apply a parameter-shift rule [24, 25] adapted for fidelity overlaps, following the formulation introduced in [26]. This method uses central finite differences with shifts of ±$\pm\pi/2$, yielding an exact expression for the metric from four circuit evaluations.
@@
where $\mathbf{e}_{j_{1}}$ and $\mathbf{e}_{j_{2}}$ are unit vectors in the parameter space. If the metric tensor is assumed to be diagonal (i.e., $g_{j_{1}j_{2}}=0$ for $j_{1}\neq j_{2}$), the expression simplifies considerably and reduces the number of required evaluations:
@@
In this scenario, the Christoffel symbols can be computed directly from the derivatives of the metric tensor using the higher-order parameter-shift rule, enabling their estimation directly from the quantum circuit. For the diagonal case, the results are summarized in the following proposition:
@@
$\Gamma^{i}_{jk}$ $=\frac{1}{16g_{ii}(\theta)}\Bigg{[}\delta_{ij}\Bigg{(}-\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}+\dfrac{\pi}{2}e_{k})\rangle\right|^{2}+\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}-\dfrac{\pi}{2}e_{k})\rangle\right|^{2}\Bigg{)}$ (29)
$+\delta_{ik}\Bigg{(}-\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}+\dfrac{\pi}{2}e_{j})\rangle\right|^{2}+\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}-\dfrac{\pi}{2}e_{j})\rangle\right|^{2}\Bigg{)}$
$-\delta_{jk}\Bigg{(}-\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{j}+\dfrac{\pi}{2}e_{i})\rangle\right|^{2}+\left|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{j}-\dfrac{\pi}{2}e_{i})\rangle\right|^{2}\Bigg{)}\Bigg{]}.$
@@
The computation of the metric tensor and Christoffel symbols involves evaluating combinations of state overlaps at each step of the update rule algorithm (24). Specifically, the Christoffel symbols (29) require six distinct expectation values. For instance, one of these overlaps can be expressed as:
$|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\pi e_{i}+\frac{\pi}{2}e_{k})\rangle|^{2}=|\langle 0|U^{\dagger}(\bm{\theta})U(\bm{\theta}+\pi e_{i}+\frac{\pi}{2}e_{k})|0\rangle|^{2}.$ (30)
This overlap can be intuitively interpreted as the probability of measuring all qubits of the quantum state $U(\bm{\theta})U(\bm{\theta}+\pi e_{i}+\frac{\pi}{2}e_{k})|0\rangle$ in the computational basis at $|0\rangle$. If $l$ denotes the number of variational parameters in the quantum circuit generating the state $|\psi(\bm{\theta})\rangle$, then computing the full Fubini–Study metric requires $O(l^{2})$ quantum circuits of depth $2d$, and the Christoffel symbols require $O(l^{3})$ quantum circuits of the same depth. Here, $d$ refers to the depth of a single execution of the variational ansatz, i.e., the number of sequential gate layers. For the diagonal approximation of the metric, the cost reduces to $O(l)$ quantum circuits, while the Christoffel symbols require $O(l^{2})$ quantum circuits, both still with depth $2d$.
VI Application Examples
@@
VI.1 Example 1 : Single Qubit
@@
where $\bm{\theta}=(\theta_{0},\theta_{1})$.
The Hamiltonian for which we want to find the ground state is $H=\sigma_{x}$, where the energy is expressed in dimensionless units in this pedagogical VQE example. Thus, the cost function for the system is:
@@
The Christoffel symbols of this metric, calculated using Eq. (4), have the following nonzero components:
@@
Figure 1: Performance of GD, QNG, and QNGGC with varying correction parameters ($b=\eta^{2}$, $b=0.05$, and $b=0.2$). The initial parameters $(\theta_{0},\theta_{1})$ are set to $[\pi/12,\pi/12]$, corresponding to a specific initial quantum state defined in Eq. (31). A learning rate $\eta=0.05$ is used for all optimizers. (a) Shows the evolution of the energy cost function over 30 iterations, demonstrating the improved convergence behavior of QNGGC relative to QNG and GD. (b) Depicts the logarithmic energy difference, $\log_{10}(\Delta E)$, where $\Delta E$ represents the gap between the current energy and the target energy $E^{*}=-1$. The current energy is computed as the expectation value $\langle\phi(\theta)|H|\phi(\theta)\rangle$ at each step. (c) Displays the norm of the geodesic correction term across different correction parameters $b$, highlighting the impact of geodesic corrections during optimization. (d) Shows the trajectory of the parameters $\theta_{0}$ and $\theta_{1}$ throughout the optimization process, illustrating how each method navigates the parameter space. The QNGGC paths are more direct, emphasizing the benefits of incorporating geodesic corrections.
Figure 1 demonstrates the impact of geodesic corrections on optimization performance. The inclusion of geodesic corrections in QNGGC significantly enhances convergence speed (in terms of steps) compared to GD and QNG, as depicted in subplots (a) and (b), where QNGGC rapidly approaches the target energy $E^{*}=-1$. Subplot (c) shows the norm of the geodesic correction term, $\frac{b}{2}\Gamma^{i}_{lm}(\theta^{\mu}_{t})\left(g^{lj}(\theta^{\mu}_{t})\partial_{j}\mathcal{L}(\theta^{\mu}_{t})\right)\left(g^{mj}(\theta^{\mu}_{t})\partial_{j}\mathcal{L}(\theta^{\mu}_{t})\right)$, which plays a crucial role in guiding the optimization process, particularly in the initial iterations. The subsequent reduction of this norm towards zero indicates that the convergence of the gradient is being effectively managed, leading to more stable updates. Subplot (d) illustrates the optimization paths in the parameter space, where QNGGC provides a more direct and efficient trajectory compared to GD and QNG. This behavior underscores the advantage of leveraging curvature corrections, resulting in more precise and adaptive parameter updates.
VI.2 Example 2: Two-Qubit Simulation of the Hydrogen Molecule ($\text{H}_{2}$)
@@
where $\alpha=0.4$ and $\beta=0.2$. In this Hamiltonian, $\sigma_{z}$ and $\sigma_{x}$ are Pauli matrices that represent spin operators acting on individual qubits. The energy is expressed in Hartree atomic units, which is standard for molecular electronic structure simulations. This Hamiltonian has four eigenvalues:
@@
In this example, we consider a variational ansatz with three parameters $\bm{\theta}=(\theta_{0},\theta_{1},\theta_{2})$, resulting in a $3\times 3$ Fubini-Study metric $F$, which is calculated as:
@@
Figure 3: Comparison of the convergence of GD, QNG, and QNGGC. The initial parameters $(\theta_{0},\theta_{1},\theta_{2})$ are set to $[-0.2,-0.2,0]$, corresponding to the initial state defined by the ansatz in Eq. (39). The learning rate is $\eta=0.05$ for all optimizers, and $b=\eta^{2}$ for QNGGC. (a) Shows energy convergence. (b) Plots $\log_{10}(\Delta E)$, where $\Delta E$ is the energy difference from the target energy $E^{*}\approx-0.82462$ Hartree. (c) Displays the norm of the geodesic correction term in QNGGC. (d) Illustrates the fidelity convergence rate. The optimization runs for 30 iterations with a regularization term $\lambda=10^{-6}$.
@@
Figure 4: Same conditions as in Figure 3, except the initial parameters $(\theta_{0},\theta_{1},\theta_{2})$ are set to $[\pi/2,\pi/2,0]$, and $b=0.4$. This configuration is more challenging due to the proximity of the initial state to a flat region of the cost landscape (plateau), which slows down optimization.
Figure 4 highlights the performance of the optimizers when initialized near a plateau in the cost landscape. Despite the flat region, QNGGC efficiently escapes and converges faster to the ground state compared to GD and QNG.
@@
VI.3 Example 3: Transverse Field Ising Model
@@
where $Z_{i}$ and $X_{i}$ are Pauli operators acting on the $i$-th qubit, with the transverse field strength denoted by $h$ and fixed at $h=10$ for our simulations. All simulations are performed in dimensionless units by setting the spin coupling strength to $J=1$; hence, both the energy and transverse field $h$ are expressed in units of $J$. This model captures the interplay between nearest-neighbor spin interactions and a transverse magnetic field, driving the system between distinct quantum phases. For smaller values of the transverse field $h$, the spins tend to align along the z-direction, corresponding to a ferromagnetic ordered phase. As $h$ increases, the system becomes increasingly disordered and transitions into a paramagnetic phase. The critical point at $h=1$ marks the quantum phase transition between these phases, where spontaneous symmetry breaking and critical quantum fluctuations occur.
@@ VII Conclusions and Outlook
While QNGGC offers improved convergence through geodesic corrections, it also introduces additional computational overhead. Gradient descent (GD) requires only $O(l)$ circuit evaluations per iteration (with $l$ denoting the number of variational parameters) and no metric-related overhead, but typically suffers from slow convergence. QNG improves convergence at a cost of $O(l^{2})$ circuits for evaluating the Fubini–Study metric. QNGGC further enhances convergence by incorporating curvature corrections via Christoffel symbols, increasing the cost to $O(l^{3})$. Although this complexity is manageable for small, shallow quantum circuits, it may become prohibitive in larger or noisier quantum circuits.
A promising direction for reducing this cost to a constant per iteration is the use of simultaneous perturbation stochastic approximation (SPSA) [36], which is particularly well suited to such noisy, deep circuits, and has already demonstrated advantages for efficiently computing the Fubini–Study metric [11]; a similar strategy could be applied to the computation of Christoffel symbols as well. In this work, we focused on small, shallow quantum circuits and the geometric structure underlying the rederivation of QNG from the Riemannian exponential map, as well as its extension to higher-order integrators that lead to QNGGC. We leave the exploration of deeper, noisy quantum circuits to future research. Once the computational cost is reduced through SPSA or other stochastic techniques, the framework could be extended further and find broader applications, such as time-dependent optimization methods inspired by differential geometry and the metric tensor formalism.
VIII Acknowledgements
@@ A.2 Fubini-Study Metric
To derive the Fubini-Study metric tensor $g_{ij}(\bm{\theta})$, we begin by examining the infinitesimal distance between nearby quantum states in the parameter space. The metric tensor captures the local geometric structure of this space, reflecting how the quantum state $|\psi(\bm{\theta})\rangle$ evolves under small displacements [6, 8, 32].
@@
Differentiating this condition with respect to the parameters $\theta^{i}$ gives:
$\left\langle\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\right\rangle+\left\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle=0.$ (45)
Differentiating Eq. (45) with respect to the parameters $\theta^{j}$, we obtain:
$\left\langle\psi(\bm{\theta})|\partial_{i}\partial_{j}\psi(\bm{\theta})\right\rangle+\left\langle\partial_{i}\partial_{j}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle+\left\langle\partial_{i}\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle+\left\langle\partial_{j}\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\right\rangle=0.$ (46)
Next, consider the quantum state $|\psi(\bm{\theta}+\bm{\delta\theta})\rangle$ near $|\psi(\bm{\theta})\rangle$, where $\bm{\delta\theta}$ is the displacement vector. Using a Taylor expansion, we express the shifted state as:
@@
$\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\bm{\delta\theta})\rangle=1+\langle\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\rangle\delta\theta^{i}+\frac{1}{2}\langle\psi(\bm{\theta})|\partial_{i}\partial_{j}\psi(\bm{\theta})\rangle\delta\theta^{i}\delta\theta^{j}+\mathcal{O}(\delta\theta^{3}).$ (48)
The fidelity between the states $|\psi(\bm{\theta})\rangle$ and $|\psi(\bm{\theta}+\delta\theta)\rangle$ is given to quadratic order in $\bm{\delta\theta}$ by:
$|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\bm{\delta\theta})\rangle|^{2}$ $=\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\bm{\delta\theta})\rangle\langle\psi(\bm{\theta}+\bm{\delta\theta})|\psi(\bm{\theta})\rangle$
$=1+\left[\left\langle\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\right\rangle+\left\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle\right]\delta\theta^{i}$
$\quad+\left[\left\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle\left\langle\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle\right]\delta\theta^{i}\delta\theta^{j}$
$\quad+\frac{1}{2}\left[\left\langle\psi(\bm{\theta})|\partial_{i}\partial_{j}\psi(\bm{\theta})\right\rangle+\left\langle\partial_{i}\partial_{j}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle\right]\delta\theta^{i}\delta\theta^{j}+\cdots$
Using Eqs. (45) and (46), we get
$|\langle\psi(\bm{\theta})|\psi(\bm{\theta}+\bm{\delta\theta})\rangle|^{2}$ $=1+\bigg{[}\left\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle\left\langle\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle$
$\quad-\frac{1}{2}\left(\left\langle\partial_{i}\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle+\left\langle\partial_{j}\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\right\rangle\right)\bigg{]}\delta\theta^{i}\delta\theta^{j}+\cdots$ (50)
Using the approximation for the Fubini-Study distance between quantum states,
$d^{2}(P_{\psi},P_{\phi})=\arccos^{2}(|\langle\psi|\phi\rangle|)\approx 1-|\langle\psi|\phi\rangle|^{2}+\mathcal{O}((1-|\langle\psi|\phi\rangle|^{2})^{2}),$ (51)
and the fidelity expansion up to second order, it follows that the infinitesimal squared distance is given by
$d^{2}(P_{\psi(\bm{\theta})},P_{\psi(\bm{\theta}+\bm{\delta\theta})})$ $=\left[\frac{1}{2}\left(\left\langle\partial_{i}\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle+\left\langle\partial_{j}\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\right\rangle\right)\right.$
$\quad\left.-\left\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle\left\langle\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle\right]d\theta^{i}d\theta^{j}.$ (52)
The first term in (52) is manifestly real. The second term is also real because from the normalization condition (as shown in Eq. (45)), we have
$\text{Re}\left[\left\langle\psi(\bm{\theta})|\partial_{i}\psi(\bm{\theta})\right\rangle\right]=0.$ (53)
Therefore, the infinitesimal squared distance is given by the real part of the quantum geometric tensor:
$d^{2}(P_{\psi(\bm{\theta})},P_{\psi(\bm{\theta}+\bm{\delta\theta})})$ $=\text{Re}\left[\left\langle\partial_{i}\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle-\left\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle\left\langle\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle\right]d\theta^{i}d\theta^{j}.$ (54)
This directly defines the Fubini-Study metric as:
$g_{ij}(\bm{\theta})=\text{Re}\left[\left\langle\partial_{i}\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle-\left\langle\partial_{i}\psi(\bm{\theta})|\psi(\bm{\theta})\right\rangle\left\langle\psi(\bm{\theta})|\partial_{j}\psi(\bm{\theta})\right\rangle\right].$ (55)